# Spirals and Beyond: Competitive Plane Search with Multi-Speed Agents

### v1

@@ 6 Conic Complement Search with a Unit-Speed Agent
To describe the trajectory for searching $\mathcal{W}_{\phi}$, we first develop some intuition for why this can be done more efficiently than searching the entire plane $\mathcal{P}$ with a unit-speed agent, even when $\phi\in(0,\pi)$ is small. A key aspect in analyzing Algorithm 1 for $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$ is that the search cost is independent of the target's angular coordinate. For this reason, and without loss of generality, we fix a reference ray $\mathcal{L}_{0}$ and perform worst-case analysis under the assumption that the target lies on $\mathcal{L}_{0}$.
Let $f_{0}^{k}(t)$ denote the trajectory of Algorithm 1, applicable to $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$. This trajectory intersects $\mathcal{L}_{0}$ at the points $X_{s}=f_{0}^{k}(2\pi s)$ and intersects the ray $\mathcal{L}_{\phi}$ at the points $Z_{s}=f_{0}^{k}(2\pi s+\phi)$, where $s\in\mathbb{Z}$, see also Figure 4. To adapt $f_{0}^{k}(t)$ for searching $\mathcal{W}_{\phi}$ (i.e., the plane excluding the cone $\mathcal{C}_{\phi}$), we can modify the trajectory by replacing the curved segment between $X_{s}$ and $Z_{s}$ with the shorter straight line segment $X_{s}Z_{s}$, while preserving the original trajectory between $Z_{s}$ and $X_{s+1}$. With this modification, every point in $\mathcal{W}_{\phi}$ is reached strictly faster. However, the search cost becomes dependent on the target's location. Since the search cost is defined as the supremum over all possible target positions in the unbounded domain $\mathcal{W}_{\phi}$, this modification does not necessarily reduce the worst-case cost.
@@
It follows that for each $s\in\mathbb{Z}$, the set of targets exposed by $p_{s}(t)$, for $t\in(\phi,2\pi)$, forms a nested family in $s$, and their union covers $\mathcal{W}_{\phi}$. Define the wedge $W_{s}\subseteq\mathcal{W}_{\phi}$ as the set of targets first exposed by $p_{s+1}(t)$ for some $t\in(\phi,2\pi)$. The wedges $\{W_{s}\}_{s}$ are disjoint, and by Lemma 6.1, we have $\bigcup_{s}W_{s}=\mathcal{W}_{\phi}$.
@@
Proof.
By Lemma 3.2, and since $\alpha>0$,
@@
using the identity $h_{s}(t)=2\pi ks+h_{0}(t)$. Also, in Cartesian coordinates, $X_{s}=e^{2\pi ks}(1,0)$ and $Y_{s}=\lambda e^{2\pi ks}(\cos\left(\phi\right),\sin\left(\phi\right))$, so
@@
Proof.
We have $p_{s+1}(\psi)\in\mathcal{L}_{\psi}$ with $\left\lVert p_{s+1}(\psi)\right\rVert=e^{h_{s+1}(\psi)}=e^{2\pi k}e^{h_{s}(\psi)},$ and by definition of $W_{s}$, this is at least $\left\lVert A\right\rVert$ for all $A\in W_{s}\cap\mathcal{L}_{\psi}$. By Lemma 6.2, the time to reach $p_{s+1}(\psi)$ is
@@
We now show that we can choose $k$ and $\lambda$ to make the search cost independent of the target's location.
Lemma 6.4.
@@
Proof.
Since $\lambda>1$, it follows that $k_{\lambda}>\frac{\ln\lambda}{2\pi}>0$, so $2\pi k_{\lambda}>\ln\lambda$, and hence $\lambda<e^{2\pi k_{\lambda}}$, satisfying the algorithm's conditions.
Let $\psi\in(\phi,2\pi)$ and $A\in W_{s}\cap\mathcal{L}_{\psi}$. By Lemma 6.3, $A$ is exposed by $p_{s+1}(\psi)$, and since
@@
We are now ready to prove Theorem 2.5, which follows from Lemma 6.4 by showing that the search cost admits a minimum for every $\phi\in(0,\pi)$ when the parameter $\lambda$ is restricted to values strictly greater than $1$.
Proof of Theorem 2.5.
By Lemma 6.4, for all $\lambda>1$, and setting $k=k_{\lambda}$, the search cost of Algorithm 5 is
@@
First, observe that $p(\lambda)$ is continuous for all $\lambda>1$, and since $\alpha_{\lambda}>0$, we have $p(\lambda)\geq 0$. Substituting the expression for $\alpha_{\lambda}$ from Lemma 6.4,
@@ 7 Sub-Optimality of the Spiral-Type Off-Set Algorithm
In this section, we demonstrate that the best performance achieved by the logarithmic-spiral-type Off-Set Algorithm 2 for planar search, as described in Theorem 2.2, can be strictly improved when the speed of one of the agents is sufficiently small. We illustrate this by focusing on the special case of two agents searching the plane, that is, the problem $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$, where $c\in(0,1)$. This establishes Theorems 2.6 and 2.7. We begin with an observation.
Proposition 7.1.
@@
Proposition 7.1 is a negative result, as it shows that slow agents with speed $c<1/\mathcal{U}_{1}$ cannot help reduce the optimal search cost below $\mathcal{U}_{1}$ in the problem $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$. However, can the Off-Set logarithmic-spiral Algorithm 2 make use of such slow-moving agents? The following immediate corollary of Theorem 2.2 addresses this.
Corollary 7.2.
@@
 Sequence $\{q_{i}\}_{i\in\mathbb{Z}}$ defined as in Lemma 5.2, and fixed to $q_{i}=q_{i}(\phi)$ by Theorem 2.4.
 Parameters $\lambda,k_{\lambda}$ defined in Lemma 6.4, and fixed to $\lambda=\lambda(\phi)$, $k_{\lambda}=k_{\lambda}(\phi)$ by Theorem 2.5.
@@
$\frac{\sin\left(\phi\right)}{2\sin\left(\phi/2\right)}\cdot\frac{\mathcal{F}_{\phi}}{\gamma(\phi)},$
@@
Theorem 2.7 now follows as a numerical corollary of Theorem 2.6. For it's proof, we numerically compute the pairs $(\gamma(\phi),\mathcal{R}_{\phi})$ for $\phi\in(0,\pi/2)$, where a speed-$\gamma(\phi)$ agent searches the cone $\mathcal{C}_{\phi}$, incurring a corresponding search cost $\mathcal{R}_{\phi}$. We compare these values to the upper bound $\mathcal{U}_{2}/\sqrt{c}$ given in Theorem 2.2, as shown in Figure 1. Since our goal is to optimize performance for small values of $c$, the approach is valid as long as the range of $\gamma(\phi)$ contains the interval
@@
Figure 6: Plot of $\gamma(\phi)=\tfrac{\sin\left(\phi\right)}{2\sin\left(\phi/2\right)}\tfrac{\mathcal{F}_{\phi}}{\mathcal{R}_{\phi}}$ against $\phi$, where $\mathcal{F}_{\phi}$ and $\mathcal{R}_{\phi}$ are defined in Theorems 2.4 and 2.5, respectively.
@@ 8 Discussion
We studied point search in the plane using multi-speed agents following spiral-based trajectories, extending a classical result for a single unit-speed agent. While spirals are known to be optimal for one agent, our results suggest that in the multi-agent setting they may become suboptimal when some agents are too slow relative to the geometric mean of all speeds. Our main result is a general upper bound on search cost that scales with the cost for unit-speed agents and is inversely proportional to the geometric mean of selected speeds. This highlights a trade-off between using all available agents and excluding slower ones to maintain efficiency. Several questions remain open. Can spiral trajectories be proven optimal, or can lower bounds be established, for some classes of speed distributions? Does a similar trade-off arise in shoreline search with heterogeneous agents? What if agent speeds are unknown or change over time? How do these strategies extend to bounded domains or environments with obstacles? We hope this work motivates further study on optimal strategies and limits of collaborative search in geometric settings.

### v2

@@ 6 Conic Complement Search with a Unit-Speed Agent
To describe the trajectory for searching $\mathcal{W}_{\phi}$, we first develop some intuition for why this can be done more efficiently than searching the entire plane $\mathcal{P}$ with a unit-speed agent, even when $\phi\in(0,\pi)$ is small. A key aspect in analyzing Algorithm 1 for $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$ is that the search cost is independent of the target’s angular coordinate. For this reason, and without loss of generality, we fix a reference ray $\mathcal{L}_{0}$ and perform worst-case analysis under the assumption that the target lies on $\mathcal{L}_{0}$.
Let $f_{0}^{k}(t)$ denote the trajectory of Algorithm 1, applicable to $\mathcal{P}\text{-}\textsc{MSP}_{1}(1)$. This trajectory intersects $\mathcal{L}_{0}$ at the points $X_{s}=f_{0}^{k}(2\pi s)$ and intersects the ray $\mathcal{L}_{\phi}$ at the points $Z_{s}=f_{0}^{k}(2\pi s+\phi)$, where $s\in\mathbb{Z}$, see also Figure 5. To adapt $f_{0}^{k}(t)$ for searching $\mathcal{W}_{\phi}$ (i.e., the plane excluding the cone $\mathcal{C}_{\phi}$), we can modify the trajectory by replacing the curved segment between $X_{s}$ and $Z_{s}$ with the shorter straight line segment $X_{s}Z_{s}$, while preserving the original trajectory between $Z_{s}$ and $X_{s+1}$. With this modification, every point in $\mathcal{W}_{\phi}$ is reached strictly faster. However, the search cost becomes dependent on the target’s location. Since the search cost is defined as the supremum over all possible target positions in the unbounded domain $\mathcal{W}_{\phi}$, this modification does not necessarily reduce the worst-case cost.
@@
It follows that for each $s\in\mathbb{Z}$, the set of targets exposed by $p_{s}(t)$, for $t\in(\phi,2\pi)$, forms a nested family in $s$, and their union covers $\mathcal{W}_{\phi}$. Define the wedge $W_{s}\subseteq\mathcal{W}_{\phi}$ as the set of targets first exposed by $p_{s+1}(t)$ for some $t\in(\phi,2\pi)$. The wedges $\{W_{s}\}_{s}$ are disjoint, and by Lemma 12, we have $\bigcup_{s}W_{s}=\mathcal{W}_{\phi}$.
@@
Proof
By Lemma 2, and since $\alpha>0$,
@@
using the identity $h_{s}(t)=2\pi ks+h_{0}(t)$. Also, in Cartesian coordinates, $X_{s}=e^{2\pi ks}(1,0)$ and
$Y_{s}=\lambda e^{2\pi ks}(\cos\left(\phi\right),\sin\left(\phi\right)),$
so
@@
Proof
We have $p_{s+1}(\psi)\in\mathcal{L}_{\psi}$ with $\left\lVert p_{s+1}(\psi)\right\rVert=e^{h_{s+1}(\psi)}=e^{2\pi k}e^{h_{s}(\psi)},$ and by definition of $W_{s}$, this is at least $\left\lVert A\right\rVert$ for all $A\in W_{s}\cap\mathcal{L}_{\psi}$. By Lemma 13, the time to reach $p_{s+1}(\psi)$ is
@@
We now show that we can choose $k$ and $\lambda$ to make the search cost independent of the target’s location.
Lemma 15
@@
Proof
Since $\lambda>1$, it follows that $k_{\lambda}>\frac{\ln\lambda}{2\pi}>0$, so $2\pi k_{\lambda}>\ln\lambda$, and hence $\lambda<e^{2\pi k_{\lambda}}$, satisfying the algorithm’s conditions.
Let $\psi\in(\phi,2\pi)$ and $A\in W_{s}\cap\mathcal{L}_{\psi}$. By Lemma 14, $A$ is exposed by $p_{s+1}(\psi)$, and since
@@
We are now ready to prove Theorem 2.4, which follows from Lemma 15 by showing that the search cost admits a minimum for every $\phi\in(0,\pi)$ when the parameter $\lambda$ is restricted to values strictly greater than $1$.
Proof(of Theorem 2.4)
By Lemma 15, for all $\lambda>1$, and setting $k=k_{\lambda}$, the search cost of Algorithm 5 is
@@
First, observe that $p(\lambda)$ is continuous for all $\lambda>1$, and since $\alpha_{\lambda}>0$, we have $p(\lambda)\geq 0$. Substituting the expression for $\alpha_{\lambda}$ from Lemma 15,
@@ 7 Sub-Optimality of the Spiral-Type Off-Set Algorithm
In this section, we demonstrate that the best performance achieved by the logarithmic-spiral-type Off-Set Algorithm 2 for planar search, as described in Theorem 2.2, can be strictly improved when the speed of one of the agents is sufficiently small. We illustrate this by focusing on the special case of two agents searching the plane, that is, the problem $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$, where $c\in(0,1)$. This establishes Theorems 2.5 and 2.6. We begin with an observation.
Proposition 2
@@
Proposition 2 is a negative result, as it shows that slow agents with speed $c<1/\mathcal{U}_{1}$ cannot help reduce the optimal search cost below $\mathcal{U}_{1}$ in the problem $\mathcal{P}\text{-}\textsc{MSP}_{2}(1,c)$. However, can the Off-Set logarithmic-spiral Algorithm 2 make use of such slow-moving agents? The following immediate corollary of Theorem 2.2 addresses this.
Corollary 2
@@
The Off-Set Algorithm 2 is spiral based in the sense that both agents are tasked with covering all directions, and the analysis enforces a shared worst case exposure across agents. In the regime of Corollary 2, this forces a dichotomy: either the slow agent does not help at all, or it participates but drives the bound to the spiral expression $\mathcal{U}_{2}/\sqrt{c}$. The hybrid strategy below avoids this global coupling by partitioning the plane into two geometrically different regions, and assigning each agent only the region it can search efficiently. The speed-$c$ agent is restricted to a cone, where directional search is effective, while the unit-speed agent covers the conic complement using the wedge search bound, and this separation is what enables an improvement over two spirals for an intermediate range of speeds.
@@
 Sequence $\{q_{i}\}_{i\in\mathbb{Z}}$ defined as in Lemma 11, and fixed to $q_{i}=q_{i}(\phi)$ by Theorem 2.3.
 Parameters $\lambda,k_{\lambda}$ defined in Lemma 15, and fixed to $\lambda=\lambda(\phi)$, $k_{\lambda}=k_{\lambda}(\phi)$ by Theorem 2.4.
@@
$\cos\left(\phi/2\right)\cdot\frac{\mathcal{F}_{\phi}}{\gamma(\phi)},$
@@
Theorem 2.6 now follows as a numerical corollary of Theorem 2.5. For it’s proof, we numerically compute the pairs $(\gamma(\phi),\mathcal{R}_{\phi})$ for $\phi\in(0,\pi/2)$, where a speed-$\gamma(\phi)$ agent searches the cone $\mathcal{C}_{\phi}$, incurring a corresponding search cost $\mathcal{R}_{\phi}$. We compare these values to the upper bound $\mathcal{U}_{2}/\sqrt{c}$ given in Theorem 2.2, as shown in Figure 1. Since our goal is to optimize performance for small values of $c$, the approach is valid as long as the range of $\gamma(\phi)$ contains the interval
@@
Figure 7: Plot of $\gamma(\phi)=\cos\left(\phi/2\right)\tfrac{\mathcal{F}_{\phi}}{\mathcal{R}_{\phi}}$ against $\phi$, where $\mathcal{F}_{\phi}$ and $\mathcal{R}_{\phi}$ are defined in Theorems 2.3 and 2.4, respectively.
@@ 8 Discussion
We studied point search in the plane using multi-speed agents, focusing on extensions of spiral-based strategies that are known to be optimal in the single-agent, unit-speed setting. While spirals provide a natural baseline, our results show that they need not be optimal once agents have heterogeneous speeds. In particular, we identify a concrete regime in which spiral-based offset strategies become suboptimal, providing the first rigorous evidence of a qualitative separation between single-speed and multi-speed search.
The quantitative improvement obtained by the conic-complement hybrid strategy is modest and confined to a narrow range of speed ratios. We do not claim that this hybrid is close to optimal, and we currently lack intuition about the structure of globally optimal strategies in this regime. Rather, the hybrid construction serves to demonstrate that assigning geometrically distinct tasks to agents of different speeds can outperform frameworks in which all agents are required to cover all directions.
At the level of subproblems, stronger statements may be possible. For directional (cone) search, the cone-search strategy appears to capture the essential geometric difficulty of the problem, and we conjecture that it is optimal in that setting. For $n$ agents of equal unit speed, we likewise conjecture that spiral-based strategies remain optimal, although establishing this lies outside the scope of the present work. Finally, for sufficiently large speed ratios, we conjecture that the offset spiral strategy analyzed here becomes optimal again, reflecting the diminishing influence of slower agents.
Randomized strategies define a largely orthogonal direction. Randomizing the global orientation of a deterministic trajectory removes the adversary’s ability to select a worst-case target angle, so that the expected performance coincides with the average-case performance of the corresponding deterministic strategy under a uniformly random target direction. A systematic comparison between deterministic and randomized strategies in the multi-speed planar search model remains open.
Several further questions arise naturally. Can matching lower bounds be established for restricted classes of speed distributions? Does a similar separation between spiral-based and non-spiral strategies arise in shoreline search with heterogeneous agents? How should strategies adapt when agent speeds are unknown or time-varying, or when search is conducted in bounded domains or environments with obstacles?